OpenStudy (anonymous):

Finish writing the equation of the line y=1/4x+? ____, given the points (-12, 0) and (4, 4)

OpenStudy (anonymous):

change it to y=1/4 x + b now find b by plugging point in equation

OpenStudy (anonymous):

0=1/4(-12)+b b=3

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

ur welcome